OpenStudy (anonymous):

8y-3=6y+17 sovle for y and tell me how yu got the answer in steps please

OpenStudy (anonymous):

y=10

OpenStudy (anonymous):

Begin by isolated y. (Getting all of them on the same side) So subtract the 6y. 8y-6y-3=17 Then add the 3 8y-6y=20 Combine the y's 2y=20 Divide by 2. y=10

OpenStudy (anonymous):

8y-3=6y+17 Since 6y contains the variable to solve for, move it to the left-hand side of the equation by subtracting 6y from both sides. 8y-3-6y=17 Since 8y and -6y are like terms, add -6y to 8y to get 2y. 2y-3=17 Since -3 does not contain the variable to solve for, move it to the right-hand side of the equation by adding 3 to both sides. 2y=3+17 Add 17 to 3 to get 20. 2y=20 Divide each term in the equation by 2. (2y)/(2)=(20)/(2) Simplify the left-hand side of the equation by canceling the common factors. y=(20)/(2) Simplify the right-hand side of the equation by simplifying each term. y=10

OpenStudy (anonymous):

now what are the 5 steps to 15-3n=n-1